# Rua Dr. Bento Teobaldo Ferraz 271, 01140-070, São Paulo - SP, Brasil

### v1

@@ Appendix A Six-dimensional Pauli matrices
The Lorentz group SO(1,5) is locally isomorphic to SU(4) and, under this identification, spinors of SO(1,5) transform as $\boldsymbol{4}$’s or $\boldsymbol{4}^{\prime}$’s of SU(4). By definition, Weyl spinors transform as a $\boldsymbol{4}$ and are denoted by an upper lower case greek index ranging from 1 to 4. Anti-Weyl spinors transform as a $\boldsymbol{4}^{\prime}$ and are denoted by a down lower case greek index ranging from 1 to 4. All other representations of SO(1,5) can be built from tensor products of $\boldsymbol{4}$’s and $\boldsymbol{4}^{\prime}$’s. The following tensor products are of particular importance
@@ Appendix C Twisted small $\mathcal{N}=4$ SCA
With the addition of $\{\widetilde{G}^{\pm},J^{\pm\pm}\}$, the above generators obey a twisted small $\mathcal{N}=4$ SCA with central charge $c$. The remaining OPE’s of this algebra are given by

### v4

@@ Appendix A Six-dimensional Pauli matrices
The Lorentz group $\rm SO(1,5)$ is locally isomorphic to SU(4) and, under this identification, spinors of $\rm SO(1,5)$ transform as $\boldsymbol{4}$’s or $\boldsymbol{4}^{\prime}$’s of SU(4). By definition, Weyl spinors transform as a $\boldsymbol{4}$ and are denoted by an upper lower case greek index ranging from 1 to 4. Anti-Weyl spinors transform as a $\boldsymbol{4}^{\prime}$ and are denoted by a down lower case greek index ranging from 1 to 4. All other representations of $\rm SO(1,5)$ can be built from tensor products of $\boldsymbol{4}$’s and $\boldsymbol{4}^{\prime}$’s. The following tensor products are of particular importance
@@ Appendix C Twisted small $\mathcal{N}=4$ SCA
With the addition of $\{\widetilde{G}^{\pm},J^{\pm\pm}\}$, the above generators obey a twisted small $\mathcal{N}=4$ SCA with central charge $c$. The remaining OPEs of this algebra are given by